OpenStudy (anonymous):

f(x) = 3x - 7 g(x) = x2 + 3x - 4 6. Find (f ◦ g)(x) 7. Find (g ◦ f(x) how do I do thos?

OpenStudy (anonymous):

work from the inside out \[f\circ g(x)=f(g(x))=f(x^2+3x-4)\] and since \[f(\heartsuit)=3\heartsuit-7\]you have \[f(x^2+3x-4)=3(x^2-3x-4)-7\]

OpenStudy (anonymous):

clean up with some algebra probably

OpenStudy (anonymous):

It's really just multiplication?